OpenStudy (anonymous):

The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Jumperman

OpenStudy (anonymous):

@Whiterabbit541 @Elsa213 @RainCloud @mathmath333

OpenStudy (anonymous):

@morganchavez @mathmath333

OpenStudy (anonymous):

hello

OpenStudy (morganchavez):

58/3 .....multiply what ever # u get by 5 and then subtract that by 58 and divide by 3 n see if u get the answer u started with

OpenStudy (anonymous):

\[length = x, width = y\] \[3y+5=x\] \[3y+5+y=58\] y=6 6*3+5=23 23+23+6+6=58, y=6, 23=x